# Proactive and Reactive Constraint Programming for Stochastic Project Scheduling with Maximal Time-Lags

### v1

@@ 5.2 Tuning of the Methods
All CP models are solved with the IBM CP solver [te17] with default settings. Most of the deterministic RCPSP/max the j10-j30 instances can be solved within 60 seconds. For, ubo50 and ubo100 instances, we fixed the time limit to 600 seconds. We tune the classical SAA with multiple scenarios $\mathsf{proactive}_{\mathsf{SAA}}$ and a heuristic approach $\mathsf{proactive}_{\gamma}$. We used $\gamma=0.9$ for $\mathsf{proactive}_{\gamma}$. For $\mathsf{proactive}_{\mathsf{SAA}}$, we sampled quantiles at $\gamma\in[0.25,0.5,0.75,0.9]$ and set a time limit of 1800 seconds. The algorithm $\mathsf{reactive}$ uses offline the $\mathsf{proactive}_{\gamma}$ algorithm, for which we fixed $\gamma=0.9$. We investigated the effect of the time limit for rescheduling and set this hyperparameter to 2 seconds. The algorithm $\mathsf{stnu}$ consists of an offline procedure, which comprises building up the network and checking dynamic controllability; if it’s not DC, the method fails. We observed that that the choice of the $\gamma$-quantile for the fixed-point schedule significantly affects the ratio of DC networks. We select $\gamma=1$ for our final method $\mathsf{stnu}$, because this setting results in much more DC networks than lower quantiles.
@@ 5.3.1 Feasibility Ratio
We first analyze the ratio of solved problems per problem set and method in Table 10. For $c=1$, the feasibility ratios of $\mathsf{proactive}_{\mathsf{SAA}}$, $\mathsf{proactive}_{0.9}$, and $\mathsf{reactive}$ are similar for instance sets j10-ubo50, with $\mathsf{reactive}$ slightly lower for ubo100. The $\mathsf{stnu}$ method has the lowest feasibility rate, but the difference narrows as problem size increases. For $c=2$, the $\mathsf{stnu}$ achieves the highest feasibility ratios due to better handling of larger variances in durations.
c, Set $\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$ $\mathsf{reactive}$
1, j10 0.65 0.85 0.85 0.85
1, j20 0.65 0.76 0.76 0.76
1, j30 0.78 0.89 0.89 0.89
1, ubo50 0.77 0.86 0.86 0.86
1, ubo100 0.84 0.91 0.91 0.88
2, j10 0.63 0.63 0.64 0.63
2, j20 0.63 0.54 0.53 0.53
2, j30 0.63 0.51 0.48 0.49
2, ubo50 0.67 0.41 0.42 0.41
2, ubo100 0.79 0.35 0.36 0.33
Table 2: Feasibility ratio.
@@ 5.3.2 Results Statistical Tests
Table 1 shows a subset of the pairwise test results for the metrics solution quality, runtime offline, and runtime online. In our Technical Appendix, we provide all test results per instance set / noise level $c$ setting. The outcomes of the test results can be used to make partial orderings of the methods, distinguishing between a strong partial ordering (Wilcoxon test) and a weak ordering (proportion test).
@@ 5.3.3 Partial Ordering Visualization
In Figure 3 - 5, an arrow $A\rightarrow B$ indicates that in the majority of the settings either $A$ is consistently better than $B$ (Wilcoxon) and/or $A$ is better than $B$ a significant number of times (proportion). Due to space constraints and for clarity, we have chosen to display only the most common pattern per metric rather than all partial orderings per instance set and noise level. Consequently, the distinction between strong and weak partial orderings is omitted in these figures. We refer to the Technical Appendix for the partial orderings per setting including the distinction between a strong and weak partial ordering.
@@ 5.4 Analysis
Figure 4 shows that the $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$ have the lowest relative offline runtime and we found no significant difference between the two. The ordering of $\mathsf{stnu}$ and $\mathsf{proactive}_{\mathsf{SAA}}$ depends on the problem size (the ordering flips for larger instances). These relative orderings are also confirmed with the magnitude test on double hits. However, we assign infinitely bad offline computation time to infeasible solutions. When executing the Wilcoxon test we include all instances for which at least one of the two methods generated a feasible solution. For that reason, a flip in the partial ordering occurs for $c=2$, ubo50 and ubo100: $\mathsf{stnu}$ shows the best performance, and $\mathsf{proactive}_{0.9}$ outperforms $\mathsf{reactive}$ according to the Wilcoxon tests. This was mainly due to the higher feasibility ratio for the better methods (see Table 10) as the results from the magnitude test on double hits contradicted Wilcoxon in these cases (we did not visualize this pattern in the main paper, but we included it in the Appendix).
@@
We observe the general partial ordering for online runtime in Figure 5. The superiority of $\mathsf{proactive}_{0.9}$ and $\mathsf{proactive}_{\mathsf{SAA}}$ are expected, as these methods only require a feasibility check online. The faster online time of the $\mathsf{stnu}$ compared to the $\mathsf{reactive}$ can be explained by the fact that the $\mathsf{stnu}$ employs a polynomial real-time execution algorithm, while $\mathsf{reactive}$ calls a deterministic CP solver multiple times. These results are confirmed with a magnitude test on double hits. Again, there are a few settings ($c=2$, ubo50 and ubo100) in which the magnitude and the Wilcoxon test contradict each other: $\mathsf{stnu}$ outperforms the $\mathsf{proactive}$ methods based on the Wilcoxon test due to higher feasibility ratios, while the magnitude test on double hits shows better online runtime for $\mathsf{proactive}_{0.9}$ and $\mathsf{proactive}_{\mathsf{SAA}}$.
5.5 Reproducibility
Together with this article, we provide a Technical Appendix in which we include explanations of the statistical tests, hyperparameter tuning, tables with test results from all pairwise comparisons that led to the partial orderings in Figure 3, 4 and 5. Furthermore, we provide a zip file comprising our code that can be used to rerun all experiments and statistical tests. In the camera-ready version of this paper, this section will refer to a public GitHub repository of our code.
@@ 6 Conclusion and Future Work
This study proposes new scheduling methods for SRCPSP/max using the latest advances in CP and STNUs. We conducted an extensive, statistical analysis to compare the proposed algorithms for the problem instances and target the existing research gap of a lacking benchmarking paper for this problem.
@@
Until now, a MIP-based proactive method was considered the best method for SRCPSP/max, although partial order schedules have shown potential in earlier research. We found that proactive methods can be improved with online rescheduling, resulting in better solution quality for the method $\mathsf{reactive}$ compared to proactive approaches $\mathsf{proactive}_{\mathsf{SAA}}$ and $\mathsf{proactive}_{\gamma}$. We find that the algorithm $\mathsf{stnu}$ that uses partial order schedules outperforms the other methods on solution quality in our evaluation. Although in general, $\mathsf{proactive}_{\gamma}$ and $\mathsf{reactive}$ have better offline computation time than $\mathsf{stnu}$, and $\mathsf{proactive}_{\mathsf{SAA}}$ and $\mathsf{proactive}_{\gamma}$ have better online computation time than $\mathsf{stnu}$, the $\mathsf{stnu}$ also showed good relative runtime results due to the polynomial time STNU-related algorithms.
In future work, the same approach could be used to evaluate other scheduling problems, and we can gain more insight into how these methods perform on both well-known problems from the literature and in practical situations. Furthermore, the set of methods could even be broadened by including sequential methods such as those provided by [Hal22] and/or machine learning-based approaches such as the graph neural network approach by [TKPdGB+23] that was developed for stochastic RCPSP.
Appendix A Tuning
@@
A.1 Time limit for CP on deterministic instances
@@
For the ubo50 instances PSP1 to PSP10 and the ubo100 instances PSP1 to PSP10, the effect of the time limits of 60 seconds, 600 seconds, and 3600 seconds respectively, is presented in Table 3 and 4. We observe that the solver status does not change when increasing the time limit to 600 seconds, and only flips 1 instance from feasible to infeasible when increasing the time limit from 600 seconds to 3600 seconds. The makespan only improves significantly for the ubo100 instances for higher time limits. In the remaining experiments, we fixed the time limit to 60 seconds for solving deterministic CPs.
@@ A.4 Hyperparameters Selection
Appendix B Statistical Tests
B.1 Wilcoxon Test
@@
B.2 Z-test for Proportion (Binomial Distribution)
@@ B.3 Magnitude Test
Appendix C Results Tables
Please find the results of the statistical test in this section. This data led to the Figures that are included in our main paper.
  • Tables 12 for noise level $c=1$ and 14 for noise level $c=2$: including the results of the Wilcoxon and proportion tests on solution quality.
  • Tables 13 for noise level $c=1$ and 15 for noise level $c=2$: including the results of the magnitude test (independent t-test) on solution quality.
  • Tables 16 for noise level $c=1$ and 18 for noise level $c=2$: including the results of the Wilcoxon and proportion tests on time offline.
  • Tables 17 for noise level $c=1$ and 19 for noise level $c=2$: including the results of the magnitude test (independent t-test) on time offline.
  • Tables 20 and Tables 22: including the results of the Wilcoxon and proportion tests on time online.
  • Tables 21 and 23: including the results of the magnitude test (independent t-test) on time online.
C.1 Feasibility Ratio
First, we analyze the feasibility ratio obtained by the different methods. For $c=1$, we observe the feasibility ratio obtained by $\mathsf{proactive}_{\mathsf{SAA}}$, $\mathsf{proactive}_{0.9}$, $\mathsf{reactive}$ are the same for instance sets j10 - ubo50, and only for ubo100 $\mathsf{reactive}$ is a little bit lower than the $\mathsf{proactive}_{\mathsf{SAA}}$, $\mathsf{proactive}_{0.9}$. The $\mathsf{stnu}$ method has the lowest feasibility rate but remarkably the difference becomes smaller when the size of the problem grows (the difference is the smallest for ubo100).
We observe a different pattern for $c=2$, with a higher noise factor. The highest feasibility ratios are obtained by $\mathsf{stnu}$. This could be explained by the fact that the $\mathsf{stnu}$ method uses the information about the distribution and is, therefore, better at handling the larger variances in activity duration.

### v2

@@ 5.2 Tuning of the Methods
All CP models are solved with the IBM CP solver [IBM24] with default settings. Most of the deterministic RCPSP/max the j10-j30 instances can be solved within 60 seconds. For, ubo50 and ubo100 instances, we fixed the time limit to 600 seconds. We tune the classical SAA with multiple scenarios $\mathsf{proactive}_{\mathsf{SAA}}$ and a heuristic approach $\mathsf{proactive}_{\gamma}$. We used $\gamma=0.9$ for $\mathsf{proactive}_{\gamma}$. For $\mathsf{proactive}_{\mathsf{SAA}}$, we sampled quantiles at $\gamma\in[0.25,0.5,0.75,0.9]$ and set a time limit of 1800 seconds. The algorithm $\mathsf{reactive}$ uses offline the $\mathsf{proactive}_{\gamma}$ algorithm, for which we fixed $\gamma=0.9$. We investigated the effect of the time limit for rescheduling and set this hyperparameter to 2 seconds. The algorithm $\mathsf{stnu}$ consists of an offline procedure, which comprises building up the network and checking dynamic controllability; if it’s not DC, the method fails. We observed that that the choice of the $\gamma$-quantile for the fixed-point schedule significantly affects the ratio of DC networks. We select $\gamma=1$ for our final method $\mathsf{stnu}$, because this setting results in much more DC networks than lower quantiles.
@@ 5.3.1 Feasibility Ratio
We first analyze the obtained feasibility ratios in Table 2. For $\epsilon=1$, the feasibility ratios of $\mathsf{proactive}_{\mathsf{SAA}}$, $\mathsf{proactive}_{0.9}$, and $\mathsf{reactive}$ are similar for instance sets j10-ubo50, with $\mathsf{reactive}$ slightly lower for ubo100. The $\mathsf{stnu}$ method has the lowest feasibility rate, but the difference narrows as problem size increases. For $\epsilon=2$, the $\mathsf{stnu}$ achieves the highest feasibility ratios due to better handling of larger variances in durations.
Set $\epsilon=1$ $\epsilon=2$
$\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$ $\mathsf{react}$ $\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$ $\mathsf{react}$
j10 0.65 0.85 0.85 0.85 0.63 0.63 0.64 0.63
j20 0.65 0.76 0.76 0.76 0.63 0.54 0.53 0.53
j30 0.78 0.89 0.89 0.89 0.63 0.51 0.48 0.49
ubo50 0.77 0.86 0.86 0.86 0.67 0.41 0.42 0.41
ubo100 0.84 0.91 0.91 0.88 0.79 0.35 0.36 0.33
Table 2: Feasibility Ratios for $\epsilon=1$ and $\epsilon=2$.
@@ 5.3.2 Results Statistical Tests
Table 1 shows a subset of the pairwise test results for the metrics solution quality, offline time, and online time. In our Appendix, we provide all test results per instance set / noise level $\epsilon$ setting. The outcomes of the test results can be used to make partial orderings of the methods, distinguishing between a strong partial ordering (Wilcoxon test) and a weak ordering (proportion test). Besides that, the results of the magnitude test give insight in the magnitude differences of each performance metric based on the double hits. For example, in Table 1, the normalized average makespan is 0.958 for $\mathsf{stnu}$ and 1.015 for $\mathsf{reactive}$, with a significant advantage for $\mathsf{stnu}$ on double hits.
@@ 5.3.3 Partial Ordering Visualization
In Figures 35, an arrow $A\rightarrow B$ indicates that in the majority of the settings either $A$ is consistently better than $B$ (Wilcoxon) and/or $A$ is better than $B$ a significant number of times (proportion). Due to space constraints and for clarity, we have chosen to display only the most common pattern per metric rather than all partial orderings per instance set and noise level. Consequently, the distinction between strong and weak partial orderings is omitted in these figures. We refer to the Technical Appendix for the partial orderings per setting including the distinction between a strong and weak partial ordering.
@@ 5.4 Analysis
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$
@@
Figure 4 shows that the $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$ have the lowest relative offline runtime and we found no significant difference between the two. The ordering of $\mathsf{stnu}$ and $\mathsf{proactive}_{\mathsf{SAA}}$ depends on the problem size (the ordering flips for larger instances). These relative orderings are confirmed with the magnitude test on double hits. However, we assign infinitely bad offline computation time to infeasible solutions. When executing the Wilcoxon test we include all instances for which at least one of the two methods generated a feasible solution. For that reason, a flip in the partial ordering occurs for $\epsilon=2$, ubo50 and ubo100: $\mathsf{stnu}$ shows the best performance, and $\mathsf{proactive}_{0.9}$ outperforms $\mathsf{reactive}$ according to the Wilcoxon tests. This was mainly due to the higher feasibility ratio for the better methods (see Table 2) as the results from the magnitude test on double hits contradicted Wilcoxon in these cases (we did not visualize this pattern in the main paper, but we included it in the Appendix).
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$smaller instanceslarger instances
@@
We observe the general partial ordering for online runtime in Figure 5. The superiority of $\mathsf{proactive}_{0.9}$ and $\mathsf{proactive}_{\mathsf{SAA}}$ are expected, as these methods only require a feasibility check online. The faster online time of the $\mathsf{stnu}$ compared to the $\mathsf{reactive}$ can be explained by the fact that the $\mathsf{stnu}$ employs a polynomial real-time execution algorithm, while $\mathsf{reactive}$ calls a deterministic CP solver multiple times. These results are confirmed with a magnitude test on double hits. Again, there are a few settings ($\epsilon=2$, ubo50 and ubo100) in which the magnitude and the Wilcoxon test contradict each other: $\mathsf{stnu}$ outperforms the $\mathsf{proactive}$ methods based on the Wilcoxon test due to higher feasibility ratios, while the magnitude test on double hits shows better online runtime for $\mathsf{proactive}_{0.9}$ and $\mathsf{proactive}_{\mathsf{SAA}}$.
$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{stnu}$$\mathsf{proactive}_{\mathsf{0.9}}$
Figure 5: Summarizing illustration of the partial ordering of the different methods for time online.
@@ 6 Conclusion and Future Work
This study introduces new scheduling methods for SRCPSP/max and statistically benchmarks them, addressing the existing research gap of a lacking benchmarking paper.
@@
Until now, proactive ($\mathsf{SORU}$-$\mathsf{H}$) methods were considered the best method for SRCPSP/max, although partial order schedules have shown potential in earlier research. We found that proactive methods can be improved with online rescheduling, resulting in better solution quality for the method $\mathsf{reactive}$ compared to proactive approaches $\mathsf{proactive}_{\mathsf{SAA}}$ and $\mathsf{proactive}_{\gamma}$. We find that the algorithm $\mathsf{stnu}$ that uses partial order schedules outperforms the other methods on solution quality in our evaluation. Although in general, $\mathsf{proactive}_{\gamma}$ and $\mathsf{reactive}$ have better offline computation time than $\mathsf{stnu}$, and $\mathsf{proactive}_{\mathsf{SAA}}$ and $\mathsf{proactive}_{\gamma}$ have better online computation time than $\mathsf{stnu}$, the $\mathsf{stnu}$ also showed good relative runtime results due to the polynomial time STNU-related algorithms.
In future work, the same approach could be used to evaluate other scheduling problems, and we can gain more insight into how these methods perform on both well-known problems from the literature and in practical situations. Temporal constraints that are defined from end-to-start can be an interesting problem domain for future work. The $\mathsf{stnu}$ method can extended in multiple directions, e.g. rearranging the order of the jobs online or using probabilistic STNs are both interesting for future work.
Furthermore, the set of methods could even be broadened by including sequential approaches [Pow22] or machine learning-based methods like the graph neural network in [TKPdGB+23] for SRCPSP.
Reproducibility Statement
All our results can be reproduced following the README from our repository.
Acknowledgements
This work is supported by the AI4b.io program, a collaboration between TU Delft and dsm-firmenich, and is fully funded by dsm-firmenich and the RVO (Rijksdienst voor Ondernemend Nederland).
Appendix A Reproducibility
We uploaded our code to make it possible to reproduce the results of our research and to facilitate others using a similar benchmarking method. Please refer to the READ.ME of our repository (github.com/kimvandenhouten/AAAI25_SRCPSPmax) to find instructions for requirements and directions to our experiment scripts, and the scripts to run the statistical tests.
Appendix B Additional Explanation Stochastic Methods
B.1 SRCPSP/max Example
Let’s reuse the example instance from the main paper (Figure 1), but now we introduce uncertainty in the durations. We now suppose that activity $d$ follows a uniform discrete distribution with $P(d=1)=0.5$ and $P(d=2)=0.5$.
B.2 STNU Method
This section gives a more comprehensive explanation of the STNU method. We showcase how the temporal network for the given problem instance looks like and how we can see whether the network is DC.
B.2.1 Step 1: Get fixed-point schedule by solving the deterministic RCPSP/max
Suppose we use $\hat{d}=\{\hat{d}_{a}=2,\hat{d}_{b}=5,\hat{d}_{c}=3,\hat{d}_{d}=2,\hat{d}_{e}=2\}$, and obtain the following fixed-point solution:
(a) Fixed-point solution with resource chains.
(b) Partial order schedule based on resource chains.
Figure 6: Partial order schedules.
B.2.2 Step 2: Construct STNU
See in Figures 7 - 9 how an STNU is constructed step by step.
start afinish astart bfinish bstart cfinish cstart dfinish dstart efinish e2-25-53-3d:1D:-21-1
Figure 7: Step 2a: For each activity two nodes are created, representing its start and end, and Step 2b: We connect the start and end of each activity with a tight link (activities $a$, $b$, $c$, $e$) or a contingent link (activity $d$).
start afinish astart bfinish bstart cfinish cstart dfinish dstart efinish e2-25-53-3d:1D:-21-1-33-2-16
Figure 8: Step 2c: The minimal and maximal time lags are modeled. These edges correspond to the precedence graph of Figure 1.
start afinish astart bfinish bstart cfinish cstart dfinish dstart efinish e2-25-53-3d:1D:-21-1-33-2-160000
[…]
B.2.3 Step 3: Check DC
The resulting network of Figure 9 does not contain any negative cycles, and therefore the network is dynamically controllable [Mor14]. Note that for this specific example, if we would have used the fixed-point solution and resource chains from Figure 2, we would have found a negative cycle in the resulting STNU and therefore no dynamic controllability, see Figure 10.
start dfinish dstart afinish astart bfinish bstart cfinish cstart efinish e2-25-53-3d:1D:-21-1-33-2-160000-7
Figure 10: Non DC network, due to negative cycle.
B.2.4 Step 4: Run RTE*
The RTE* [Pos22, HP24] algorithm provides the online scheduling strategy that transforms the network into a schedule (i.e. it assigns execution times to each node). The RTE* adapts the execution times to the outcome of the contingent time points, resulting in schedule $\{s_{a}=0,s_{b}=3,s_{c}=4,s_{d}=2,s_{e}=3\}$ if $d_{d}=1$ and $\{s_{a}=0,s_{b}=4,s_{c}=5,s_{d}=2,s_{e}=4\}$ if $d_{d}=2$.
Appendix C Tuning
@@
C.1 Comparing CP and MIP for RCPSP/max
This section shows a comparison of a CP-based formulation and MILP-based formulation of the RCPSP/max problem. The MILP model that we use in the comparison is the following:
MILP Formulation for RCPSP/max
Sets:
[…]
Parameters:
$p_{j}$ $:\text{duration of task }j,\quad\forall j\in J,$
$b_{r}$ $:\text{capacity of resource }r,\quad\forall r\in R,$
$l_{r,j}$ $:\text{demand of task }j\text{ on resource }r,\quad\forall r\in R,j\in J,$
[…]
Decision Variables:
$x_{j,t}$ :$:\begin{cases}1&\text{if task }j\text{ starts at time }t,\\ 0&\text{otherwise},\end{cases}\quad\forall j\in J,t\in T,$
[…]
The MIP model is given by:
[…]
$\sum_{t\in T}t\cdot x_{j,t}+p_{j},\leq Makespan;$ $\ \forall j\in J$
[…]
$\sum_{j\in J}\sum_{\tau=t-p_{j}+1}^{t}l_{r,j}\cdot x_{j,\tau}\leq b_{r};$ $\ \forall r\in R,\,\forall t\in T$
$\sum_{t\in T}t\cdot x_{i,t}\geq 0;$ $\ \forall j\in J$
$\sum_{t\in T}x_{j,t}=1;$ $\ \forall j\in J$
Comparison MILP and CP
Instance folder # Instances Average time CP Average of time MIP
j10 50 0,02 0,41
j20 50 24,34 17,04
j30 50 1,28 68,58
ubo50 50 75,09 210,96
ubo100 50 135,30 510,75
Table 3: Runtime CP vs MILP, timelimit set to 600 seconds.
Instance folder # Instances # Count CP # Count MIP
j10 50 0 0
j20 50 2 0
j30 50 0 5
ubo50 50 6 14
ubo100 50 11 39
Table 4: Count hit timelimit (10 minutes) CP vs MILP
C.2 Time Limit for CP on Deterministic Instances
@@
For the ubo50 instances PSP1 to PSP10 and the ubo100 instances PSP1 to PSP10, the effect of the time limits of 60 seconds, 600 seconds, and 3600 seconds respectively, is presented in Table 5 and 6. We observe that the solver status does not change when increasing the time limit to 600 seconds, and only flips 1 instance from feasible to infeasible when increasing the time limit from 600 seconds to 3600 seconds. The makespan only improves significantly for the ubo100 instances for higher time limits. In the remaining experiments, we fixed the time limit to 60 seconds for solving deterministic CPs.
@@ C.5 Hyperparameters Selection
Appendix D Statistical Tests
D.1 Wilcoxon Test
@@
D.2 Z-test for Proportion (Binomial Distribution)
@@ D.3 Magnitude Test
Appendix E Results
Please find the results of the statistical test in this section. In Section E.1 we discuss the feasibility rations. In Section E.2 we present the visualizations of the partial orderings. In Section E.3, we present the results of the magnitude tests on the double hits that can give some insight into the magnitude differences of the performance metrics between the different methods.
The test results from the Wilcoxon test and the proportion tests are used to make figures for the partial orderings, and the complete tables with the results can be found in the following tables at the end of this document.
  • Table 20 for noise level $\epsilon=1$ and Table 21 for noise level $\epsilon=2$: including the results of the Wilcoxon and proportion tests on solution quality.
  • Table 22 for noise level $\epsilon=1$ and Table 23 for noise level $\epsilon=2$: including the results of the Wilcoxon and proportion tests on time offline.
  • Table 24 for noise level $\epsilon=1$ and Table 25 for noise level $\epsilon=2$: including the results of the Wilcoxon and proportion tests on online computation time.
E.1 Feasibility Ratio
First, we analyze the feasibility ratio obtained by the different methods. For $\epsilon=1$, we observe the feasibility ratio obtained by $\mathsf{proactive}_{\mathsf{SAA}}$, $\mathsf{proactive}_{0.9}$, $\mathsf{reactive}$ are the same for instance sets j10 - ubo50, and only for ubo100 $\mathsf{reactive}$ is a little bit lower than the $\mathsf{proactive}_{\mathsf{SAA}}$, $\mathsf{proactive}_{0.9}$. The $\mathsf{stnu}$ method has the lowest feasibility rate but remarkably the difference becomes smaller when the size of the problem grows (the difference is the smallest for ubo100).
We observe a different pattern for $\epsilon=2$, with a higher noise factor. The highest feasibility ratios are obtained by $\mathsf{stnu}$. This could be explained by the fact that the $\mathsf{stnu}$ method uses the information about the distribution and is, therefore, better at handling the larger variances in activity duration.